# GIFT: Unlocking Full Potential of Labels in Distilled Dataset at Near-zero Cost

### v1

@@ 5.1 Experiment Setup
We conduct experiments on both large-scale and small-scale datasets, including the full $224\times 224$ ImageNet-1k [6], Tiny-ImageNet [17] and CIFAR-100 [16]. Following previous dataset distillation studies [36, 3, 44, 5, 11], we employ ConvNet [11] and ResNet-18 [12] as our backbone architectures across all datasets. Specifically, Conv-3 is employed for CIFAR-10/100, while Conv-4 is used for Tiny-ImageNet and ImageNet-1K. For cross-architecture experiments, we additionally utilize large-scale networks, including ResNet-101 [12] and Swin-V2-Tiny [19] and small-scale networks, such as EfficientNet-B0 [30], and MobileNet-V2 [24] to verify the generalizability of our approach.
@@
We benchmark our method GIFT against state-of-the-art dataset distillation methods that leverage soft labels and scale to large high-resolution datasets, including SRe2L [36], RDED [28], DATM [11], and G-VBSM [25]. We do not compare with TESLA [5] due to the unavailability of its synthetic dataset and its inability to surpass methods like RDED [28]. Additionally, DATM only provides synthetic datasets on ConvNet and does not provide the ImageNet-1k synthetic dataset.
Knowledge distillation [13] is a straightforward approach that utilizes both hard and soft label information. Thus, we also compare our method GIFT with state-of-the-art knowledge distillation methods, including KD [13], WSLD [45], DKD [43], and NKD [35]. Further details on these methods are provided in Appendix C.
@@
Our method does not involve any distilling datasets process. We obtain all synthetic datasets directly from the source data provided by the authors 444 $\star$ SRe2L: https://github.com/VILA-Lab/SRe2L $\star$ RDED: https://github.com/LINs-lab/RDED $\star$ DATM: https://gzyaftermath.github.io/DATM/ $\star$ G-VBSM: https://github.com/shaoshitong/G_VBSM_Dataset_Condensation and simply replace the loss function. Thus, our method is a plug-and-play approach that can be easily integrated into existing dataset distillation pipelines without necessitating additional dataset synthesis or modification. Ahe experiments are conducted using an NVIDIA RTX 4090 GPU.
@@
We provide detailed hyperparameter configurations for our distilled dataset evaluation in Appendix C. Following recent works [36, 25], the evaluation on all datasets uses the parameters outlined in Table 11. We set the coefficient of label smoothing $\alpha=0.1$ and the weight hyper-parameter $\gamma=0.1$ for all methods across various synthetic datasets, as $\gamma=0.1$ is validated to be optimal through experiments depicted in Section 5.6
Table 1: Comparison with the state-of-the-art methods of dataset distillation on CIFAR-100 and Tiny-ImageNet.In the table, “-” are absent due to scalability problems.
@@ 5.2 Can GIFT Improve Performance of Dataset Distillation?
Small-Scale Dataset Comparison
@@
Large-Scale Dataset Comparison
In the large-scale full 224 $\times$ 224 ImageNet-1k dataset, as shown in Table 2, our proposed method consistently improves all baseline methods across all IPC values in 10, 50, 100, using both ConvNet and ResNet-18 as evaluation models. Specifically, compared with the current state-of-the-art methods RDED and G-VBSM, GIFT achieves significant performance gains of 1.8%and 1.2% on ResNet-18 when IPC =10, respectively. The substantial performance improvements obtained by GIFT demonstrate its capability to effectively scale to large datasets.
@@ 5.3 Can Knowledge Distillation work?
A straightforward approach to combining soft labels and hard labels is knowledge distillation [13], which transfers knowledge from a teacher model to a student model using hard labels (cross-entropy loss) and soft labels provided by a strong teacher model (KL divergence loss). In Table 3, we compare our proposed method, GIFT, with the state-of-the-art knowledge distillation techniques across synthetic datasets distilled via RDED [28]  [^5] RDED is the current state-of-the-art method as shown in Table 1, so we conduct the experiment on its synthetic datasets.
It can be observed that GIFT outperforms all knowledge distillation methods. We attribute the failure of knowledge distillation methods to the extremely small size of distilled datasets, which significantly hampers the performance of knowledge distillation methods, as corroborated by [26]. Therefore, knowledge distillation is not well-suited for our problem, further highlighting the effectiveness of GIFT.
Table 3: Comparison with the knowledge distillation methods on the synthetic dataset via RDED [28] using ConvNet. In this table, bold means the best result, underlinedmeans the second best, and ($\uparrow$) denotes improvements over the second best baseline.
@@ 5.4 Can GIFT Improve Generalization?
Cross-Architecture Generalization
@@
The success of our method is attributed to its stability. In cross-architecture scenarios, soft labels may not be sufficient due to architectural differences [31]. However, our cosine similarity approach inherently includes a normalization operation, mitigating the negative impact of label fluctuation. These results are promising, indicating that our method, which does not incur additional computational costs, is well-suited for applications involving large-scale models.
@@
Table 5: Top-1 accuracy on cross-optimization generalization on Tiny-ImageNet and CIFAR100 when IPC=10.We evaluate the performance of synthetic datasets across various optimizers. ($\uparrow$) denotes improvements over these methods.
@@
It is notable that SRe2L and RDED perform poorly in this cross-optimizer challenge. However, with our method, performance increased by 42.6% and 51.5% using SGD on CIFAR-100. A major reason for their failure is that the KL divergence loss function they use leads to high gradient variance and instability in extremely small dataset training [9]. In contrast, AdamW maintains optimization stability across optimizers by mitigating overfitting in small datasets through explicit weight decay, enhancing generalization and preventing excessive parameter growth. Our method involves cosine similarity, which computes the angle between vectors rather than their magnitude, making it less sensitive to the scale of the input data. This leads to more stable gradients, especially in scenarios with small dataset sizes.
5.5 Application: Continual Learning
Building upon many prior studies [41, 46, 14, 36] that leverage synthetic datasets in continual learning to assess the quality of synthetic data, we employ the GDumb framework [23] for our continual learning setup. This framework sequentially stores prior training data in memory and utilizes both new and stored data to train the model.
(a) 5-step
(b) 10-step
@@ 5.6 Ablation Study
To validate the efficacy and superiority of GIFT, we conducted comparative experiments with existing loss functions on synthetic datasets, utilizing both hard and soft labels. The results, presented in Table 6, clearly demonstrate that GIFT consistently outperforms all other loss functions and their combinations.
For hard label utilization, the cross-entropy (CE) loss exhibits subpar performance, mainly due to the limited information content in typically small synthetic datasets. For soft label utilization, MSE marginally outperforms the KL divergence, consistent with observations in [15]. In summary, existing loss functions in synthetic datasets fail to fully exploit the potential of all labels. In contrast, our method leverages both hard and soft labels simultaneously, thereby maximizing label utilization potential and improving performance.
Table 6: Comparsion with loss functions employed in dataset distillation. The experiment is conducted on synthetic datasets distilled via RDED [28]. In this table, boldmeans the best result, underlinedmeans the second best, and ($\uparrow$)denotes improvements over the second best baseline.
@@
GIFT (ours) 49.5 $\pm$ 0.2($\uparrow$ 1.3) 57.0 $\pm$ 0.1($\uparrow$ 0.7) 42.9 $\pm$ 0.2($\uparrow$ 1.2) 47.5 $\pm$ 0.1($\uparrow$ 0.2) 21.7 $\pm$ 0.6($\uparrow$ 1.0) 39.5 $\pm$ 0.1($\uparrow$ 0.7) 42.5$\pm$ 0.1($\uparrow$ 0.6)
@@
We conducted an ablation study to assess the necessity of the teacher label refinement and cosine similarity loss function on the small-scale dataset in Table 7 and the large-scale dataset in Table 13 in Appendix D. This evaluation was performed on both optimization-based (SRe2L) and non-optimization-based (RDED) synthetic datasets. We compared the complete method with variants lacking either the teacher label refinement or the cosine similarity loss function. In the absence of both modules, the method was trained using its native loss function.
It is obvious that in scenarios where only one module was employed, the cosine similarity loss function significantly enhanced performance due to its direct label utilization. Teacher label refinement consistently enhances performance, regardless of the presence of cosine similarity loss function, demonstrating its effectiveness. Thus, both modules are essential for performance enhancement, consistent with our analysis in Section 4.
Table 7: Ablation study of label refinement (Refine) and cosine similarity loss function (Loss) on CIFAR 100 and Tiny-ImageNet when IPC=10. This evaluation was conducted on both optimization-based (SRe2L [36]) synthetic datasets and non-optimization-based (RDED [28]) synthetic datasets using ConvNet.
@@
We examined the impact of the weight hyperparameter $\gamma$, defined in Section 4. As shown in Figure 3, GIFT achieves optimal performance when $\gamma=0.1$. Notably, for values of $\gamma$ greater than 0.1, a significant performance decline is observed across all methods as $\gamma$ increases. A plausible explanation is that larger values of $\gamma$ diminish the intra-class information content in soft labels. This observation aligns with our findings in Table 6, where training with exclusively hard labels via CE results in poor performance. To verify that $\gamma=0.1$ is also optimal for different network architectures, we conducted experiments on the same three datasets using ConNet. The results, shown in Figure 8 in Appendix D, support the same conclusion.
@@
Figure 3: Top-1 accuracy for the state-of-the-art dataset distillation methods on various synthetic datasets when IPC =10 on ResNet-18 with different $\gamma$.
6 Conclusion and Limitation
This work introduces a novel perspective on dataset distillation by emphasizing the full utilization of synthetic labels. We first conduct a comprehensive comparison of existing loss functions used for soft labels in dataset distillation. Our findings reveal that models trained on synthetic datasets exhibit significant sensitivity to the choice of loss function. Based on these insights, we propose a simple yet effective plug-and-play method, GIFT, designed to fully exploit synthetic labels. Experimental results across various scales and resolutions of image datasets demonstrate that GIFT consistently outperforms state-of-the-art dataset distillation algorithms. In this work, we mainly focus on the image classification task. However, we believe the extensions of this work to other tasks such as text data [8] and architectures like LSTM [10].

### v2

@@ 5.1 Experiment Setup
We conduct experiments on both large-scale and small-scale datasets, including the full $224\times 224$ ImageNet-1k (Deng et al., 2009), Tiny-ImageNet (Le & Yang, 2015) and CIFAR-100 (Krizhevsky et al., 2009). Following previous dataset distillation studies (Yin et al., 2023; Cazenavette et al., 2022; Zhao et al., 2023; Cui et al., 2023; Guo et al., 2024), we employ ConvNet (Guo et al., 2024) and ResNet-18 (He et al., 2016) as our backbone architectures across all datasets. Specifically, Conv-3 is employed for CIFAR-10/100, while Conv-4 is used for Tiny-ImageNet and ImageNet-1K. For cross-architecture experiments, we additionally utilize large-scale networks, including ResNet-101 (He et al., 2016) and Swin-V2-Tiny (Liu et al., 2021) and small-scale networks, such as EfficientNet-B0 (Tan & Le, 2019), and MobileNet-V2 (Sandler et al., 2018) to verify the generalizability of our approach.
@@
We benchmark our method GIFT against state-of-the-art dataset distillation methods. We categorize current state-of-the-art methods based on two key factors: scalability to ImageNet-1K and the utilization of soft labels. These categorizations are summarized in Table 13 in Appendix C. Given that our primary focus is on enhancing the use of soft labels in dataset distillation, we restrict our comparisons to methods that leverage soft labels, including SRe2L (Yin et al., 2023), RDED (Sun et al., 2024), DATM (Guo et al., 2024), G-VBSM (Shao et al., 2024), and CDA (Yin & Shen, 2023) Additionally, DATM only provides synthetic datasets on ConvNet and does not provide the ImageNet-1k synthetic dataset. CDA provides higher IPC synthetic datasets of Tiny-ImageNet and ImageNet-1k, distilled using ResNet architectures, so we mainly compare with CDA in Section 5.2.
Knowledge distillation (Hinton et al., 2015) is a straightforward approach that utilizes both hard and soft label information. Thus, we also compare our method GIFT with state-of-the-art knowledge distillation methods, including KD (Hinton et al., 2015), WSLD (Zhou et al., 2021), DKD (Zhao et al., 2022), and NKD (Yang et al., 2023). Further details on these methods are provided in Appendix C.
@@
Our method does not involve any distilling datasets process. We obtain all synthetic datasets directly from the source data provided by the authors 555 $\star$ SRe2L: https://github.com/VILA-Lab/SRe2L $\star$ RDED: https://github.com/LINs-lab/RDED $\star$ DATM: https://gzyaftermath.github.io/DATM/ $\star$ G-VBSM: https://github.com/shaoshitong/G_VBSM_Dataset_Condensation $\star$ CDA: https://github.com/VILA-Lab/SRe2L/tree/main/CDA . Notably, distilled data is generalized using both ConvNet and ResNet-18. We replace the loss function during evaluation. Thus, our method is a plug-and-play approach that can be easily integrated into existing dataset distillation pipelines without additional dataset synthesis or modification.
For the data augmentation of synthetic datasets, only synthetic datasets generated via DATM (Guo et al., 2024) are processed using ZCA whitening, as these datasets were initially distilled through ZCA whitening. Other distilled datasets are processed using DSA (Zhao & Bilen, 2021), as detailed in Table 14. All experiments are conducted using an NVIDIA RTX 4090 GPU.
@@
We provide detailed hyperparameter configurations for our synthetic dataset evaluation in Appendix C. Following recent works (Yin et al., 2023; Shao et al., 2024), the evaluation on all datasets uses the parameters outlined in Table 15. We set the coefficient of label smoothing $\alpha=0.1$ and the weight hyper-parameter $\gamma=0.1$ for all methods across various synthetic datasets, as $\gamma=0.1$ is validated to be optimal through experiments depicted in Section 5.6 and Section 8 in Appendix D.
Table 1: Comparison with the state-of-the-art methods of dataset distillation on CIFAR-100 and Tiny-ImageNet. In this table, “-” are absent due to scalability.
@@ 5.2 Can GIFT Improve Performance of Dataset Distillation?
Small-Scale Dataset Comparison.
@@
Large-Scale Dataset Comparison.
In the large-scale ImageNet-1k dataset, as reported in Table 2, our proposed method GIFT consistently improves all baseline methods across all IPC values in 10, 50, 100, using both ConvNet and ResNet-18 as evaluation models. Specifically, compared with the current state-of-the-art methods RDED and G-VBSM, GIFT achieves significant performance gains of 1.8%and 1.2% on ResNet-18 when IPC =10, respectively. The substantial performance improvements obtained by GIFT demonstrate its capability to effectively scale to large-scale datasets.
Table 3: Comparison with CDA under higher IPC on small-scale Tiny-ImageNet and large-scale ImageNet-1K on ResNet-18. In the table, ($\uparrow$) means the improvements over CDA.
Tiny-ImageNet ImageNet-1K
Method 50 100 50 100 200
CDA 49.5 $\pm$ 0.4 53.5 $\pm$ 0.3 53.7 $\pm$ 0.3 58.3 $\pm$ 0.3 63.4 $\pm$ 0.2
CDA + Ours 54.5 $\pm$ 0.3 ($\uparrow$ 5.0) 56.6 $\pm$ 0.2 ($\uparrow$ 3.1) 54.8 $\pm$ 0.2 ($\uparrow$ 1.1) 59.0 $\pm$ 0.2 ($\uparrow$ 0.8) 63.9 $\pm$ 0.1 ($\uparrow$ 0.5)
Comparison under Higher IPC.
Given that only CDA (Yin & Shen, 2023) provides a higher IPC synthetic dataset distilled, our comparison primarily centers on CDA. The results in Table 3 demonstrate that our GIFT method substantially improves the performance of CDA. Furthermore, it also achieves significant enhancements at higher IPC.
Table 4: Comparison with RDED Large-scale Netwrok. In the table, ($\uparrow$) means the improvements over RDED.
Tiny-ImageNet ImageNet-1K
Method 10 50 10 50
RDED 47.1$\pm$ 0.3 55.1 $\pm$ 0.3 42.3 $\pm$ 0.2 58.6 $\pm$ 0.1
RDED + Ours 48.6 $\pm$ 0.2 ($\uparrow$ 1.5) 56.2 $\pm$ 0.2 ($\uparrow$ 1.1) 43.5 $\pm$ 0.2 ($\uparrow$ 1.2) 59.4 $\pm$ 0.2 ($\uparrow$ 0.8)
Comparison on Large-scale Network.
In addition to conventional networks like ResNet-18, we employ the state-of-the-art dataset distillation method, RDED, to generate distilled data for Tiny-ImageNet and ImageNet-1K using Swin Transformer (Liu et al., 2021). The results, as shown in Table 4, indicate notable enhancements, verifying the effectiveness and promise of our method.
@@ 5.3 Can Knowledge Distillation work?
A straightforward approach to combining soft labels and hard labels is knowledge distillation (Hinton et al., 2015), which transfers knowledge from a teacher model to a student model using hard labels (cross-entropy loss) and soft labels provided by a strong teacher model (KL divergence loss). In Table 5, we compare our proposed method, GIFT, with the state-of-the-art knowledge distillation techniques across synthetic datasets distilled via RDED (Sun et al., 2024)  [^6] RDED is the current state-of-the-art method as shown in Table 1, so we conduct the experiment on its synthetic datasets..
It can be observed that GIFT outperforms all knowledge distillation methods. We attribute the failure of knowledge distillation methods to the extremely small size of synthetic datasets, which significantly hampers the performance of knowledge distillation methods, as corroborated by (Stanton et al., 2021). Therefore, knowledge distillation is not well-suited for our problem, further highlighting the effectiveness of GIFT.
Table 5: Comparison with the knowledge distillation methods on the synthetic dataset via RDED (Sun et al., 2024) using ConvNet. In this table, bold means the best result, underlined means the second best, and ($\uparrow$) denotes improvements over the second best baseline.
@@
5.4 Can GIFT achieve Near-zero Cost?
We perform experiments to evaluate memory and training time costs using ResNet-18 across datasets with varying scales and resolutions, as presented in Table 6. Our method demonstrates no additional peak memory usage and incurs negligible computational overhead. This efficiency is due to its emphasis on label refinement and the implementation of a general and simple loss function during the evaluation phases. Importantly, despite the negligible additional cost, GIFT yields significant performance improvements across datasets of varying scales and resolutions.
Table 6: Training Time (s) and memory (GB) costs on three synthetic datasets using ResNet-18 when IPC=10. (+) denotes additional cost over these methods.
CIFAR-100 Tiny-ImageNet ImageNet-1K
Method Training Time Memory Training Time Memory Training Time Memory
SRe2L 180.22 0.69 1249.66 2.01 2011.24 2.32
SRe2L + Ours 181.36 (+ 1.14) 0.69 1275.75 (+ 26.09) 2.01 2078.59 (+ 67.35) 2.32
RDED 171.97 0.69 1272.28 2.01 2066.35 2.32
RDED + Ours 175.17 (+ 3.20) 0.69 1298.73 (+ 26.45) 2.01 2124.61 (+ 58.26) 2.32
DATM 132.51 1.36 1018.31 12.55 - -
DATM + Ours 139.06 (+ 6.55) 1.36 1039.65 (+ 21.34) 12.55 - -
G-VBSM 183.90 0.69 1256.35 2.01 2074.72 2.32
G-VBSM + Ours 187.83 (+ 3.93) 0.69 1282.63 (+ 26.28) 2.01 2129.98 (+ 55.26) 2.32
@@ 5.5 Can GIFT Improve Generalization?
Cross-Architecture Generalization.
@@
The success of our method is attributed to its stability. In cross-architecture scenarios, soft labels may not be sufficient due to architectural differences (Vyas et al., 2020). However, our cosine similarity approach inherently includes a normalization operation, mitigating the negative impact of label fluctuation. These results are promising, indicating that our method, which does not incur additional computational costs, is well-suited for applications involving large-scale models.
@@
Table 8: Top-1 accuracy (%) on cross-optimization generalization on Tiny-ImageNet and CIFAR100 when IPC =10.We evaluate the performance of synthetic datasets across various optimizers.
@@
It is notable that SRe2L and RDED perform poorly in this cross-optimizer challenge. However, with our method, performance increased by 42.6% and 51.5% using SGD on CIFAR-100. We present an emperical and theoretical analysis of cross-optimizer generalization in Appendix E.
@@ 5.6 Ablation Study
To validate the superiority of GIFT, we compare with existing loss functions, utilizing both hard and soft labels. The results, presented in Table 9, clearly demonstrate that GIFT consistently outperforms other loss functions and their combinations.
For hard label utilization, the cross-entropy (CE) loss exhibits subpar performance, mainly due to the limited information content in typically small synthetic datasets. For soft label utilization, Jensen-Shannon (JS) divergence loss marginally outperforms the KL divergence, consistent with observations in (Kim et al., 2021). In summary, existing loss functions in synthetic datasets fail to fully exploit the potential of all labels. In contrast, our method leverages both hard and soft labels simultaneously, thereby maximizing label utilization potential and improving performance.
Table 9: Comparsion with loss functions employed in dataset distillation. The experiment is conducted on synthetic datasets distilled via RDED (Sun et al., 2024). In this table, bold means the best result, underlined means the second best, and ($\uparrow$) denotes improvements over the second best baseline.
@@
GIFT (ours) 50.6 $\pm$ 0.3 ($\uparrow$ 2.4) 57.9 $\pm$ 0.2 ($\uparrow$ 1.6) 44.0 $\pm$ 0.2($\uparrow$ 2.3) 48.3 $\pm$ 0.1 ($\uparrow$ 1.0) 24.0 $\pm$ 0.8($\uparrow$ 3.3) 39.5 $\pm$ 0.1 ($\uparrow$ 0.7) 42.5$\pm$ 0.1 ($\uparrow$ 0.6)
@@
We conduct an ablation study to assess the necessity of the label refinement and cosine similarity loss function on the small-scale dataset in Table 10 and the large-scale dataset ImageNet-1K in Table 17 in Appendix D. We compare the complete method with variants lacking either the teacher label refinement or the cosine similarity loss function. In the absence of both modules, the method is trained using its native loss function.
It is obvious that when only one module is employed, the cosine similarity loss function significantly enhances performance due to its direct label utilization. Label refinement consistently enhances performance, regardless of the presence of cosine similarity loss function, demonstrating its effectiveness. Thus, both modules are essential for enhancement, consistent with our analysis in Section 4.
Moreover, to verfiy the efficacy of label refinement, we compare the label accuracy before and after refinement. The results, shown in Figure 10 in Appendix D, demonstrate that it leads to significant performance improvements, highlighting the critical role of the proposed label refinement.
Table 10: Ablation study of label refinement (Refine) and cosine similarity loss function (Loss) on CIFAR 100 and Tiny-ImageNet when IPC =10. This evaluation is conducted on both optimization-based (SRe2L (Yin et al., 2023)) synthetic datasets and non-optimization-based (RDED (Sun et al., 2024)) synthetic datasets.
@@
We examine the impact of the weight hyperparameter $\gamma$, defined in Section 4. As shown in Figure 2, GIFT achieves optimal performance when $\gamma=0.1$ across various datasets. This consistency is attributed to the fact that the soft labels are generated by pre-trained models. Specifically, both RDED and SRe2L utilize the same pre-trained model.
Notably, for values of $\gamma$ greater than 0.1, a significant performance decline is observed across all methods as $\gamma$ increases. A plausible explanation is that larger values of $\gamma$ diminish the intra-class information content in soft labels. This observation aligns with our findings in Table 9, where training with exclusively hard labels via CE results in poor performance. To verify that $\gamma=0.1$ is also optimal for different settings, we conduct experiments on different network architectures and augmentation, as shown in Figure 8 and Figure 9 in Appendix D.
Can GIFT Enhance utilization of Hard and Smoothed Labels?
To evaluate the efficacy of GIFT across various data types, we conduct experiments on both distilled and randomly selected datasets, employing hard and smoothed labels. The results are presented in Table 18 and Table 19 in Appendix D. Obviously, GIFT consistently enhances label utilization across various label types.
@@
Figure 2: Top-1 accuracy (%) for the state-of-the-art dataset distillation methods on various synthetic datasets when IPC =10 on ResNet-18 with different $\gamma$.
6 Conclusion
This work introduces a novel perspective on dataset distillation by emphasizing the full utilization of synthetic labels. We first conduct a comprehensive comparison of existing loss functions used for soft labels in the field of dataset distillation. Our findings reveal that models trained on synthetic datasets exhibit significant sensitivity to the choice of loss function. Building on these observations, we propose a simple yet effective plug-and-play method, GIFT, which fully leverages synthetic labels without requiring additional information. The method incorporates label refinement and introduces a cosine similarity-based loss function. Furthermore, we provide a theoretical analysis to substantiate the use of cosine similarity. Experimental results across various scales and resolutions of image datasets demonstrate that GIFT consistently achieves superior performance compared to state-of-the-art dataset distillation methods.
Acknowledgment
We thank anonymous reviewers for their precious comments and feedback. This work was supported in part by the National Science and Technology Major Project (No. 2022ZD0115101), Research Center for Industries of the Future (RCIF) at Westlake University, Westlake Education Foundation, and Westlake University Center for High-performance Computing.